OpenStudy (anonymous):

Can someone just check this to see if my answers are right, or if i'm on the right path? consider the function f(x)=5x^2-17x+6 1. is the point (-7,132) on the graph of f? my answer: 132= 5(-7)^2-17(-7)+6 132=370 answer: NO 2. find f(x) when x=8 f(8)=5(8)^2-17(8)+6 f(8)=178 3. find x when f(x)=7.75 (hint use the quadratic formula) my answer: 7.75=5x^2-17x+6 0=5x^2-17x-1.75 a=5, b=17, c=-1.75 ends up to -17+18/10 or -17-18/10 =0.1 or -3.5 does any of this look right? any help would be great. thank you!

OpenStudy (anonymous):

i think your second answer is incorrect

OpenStudy (anonymous):

am i supposed to find points on that one?

OpenStudy (anonymous):

nah i think you just made a mistake in your math, just try it again

OpenStudy (anonymous):

#2 Should be 190, the process is correct but the answer is wrong.

OpenStudy (anonymous):

f(8)=5(8)^2-17(8)+6 f(8)=320-136+6 f(8)=190 is that right?

OpenStudy (anonymous):

also your last answer has the wrong sign for 17, since the quadratic equation starts with -b

OpenStudy (anonymous):

yup 190 is correct

OpenStudy (anonymous):

oh your right!

OpenStudy (anonymous):

i mean you did apply -b, but your value for b was incorrect. it looks like you got everything now though

OpenStudy (anonymous):

but the answe is still the same bc -17 squares would be a positive number anyway right?

OpenStudy (anonymous):

yeah it doesnt change the value in the radical, but you need to look at the value before it, the -b plus or minus, b/c -b = -(-17) = 17

OpenStudy (anonymous):

again, you got the right process but, the designation of b should be -17

OpenStudy (anonymous):

oh. okay. i get it. duh! thank you

OpenStudy (anonymous):

glad you got it :) np

OpenStudy (anonymous):

therefore applying the formula \[-b \pm \sqrt{b ^{2}-4ac} all \over 2a\]

OpenStudy (anonymous):

okay here i am: 17+18/10=3.5 and 17-18/10=-0.1 i'm not getting that 1.1 you got?

OpenStudy (anonymous):

disrecard the *all please that's right. the answers are {3.5 and -0.1} only

OpenStudy (anonymous):

fabulous! thank you soo so much!

OpenStudy (anonymous):

just a typo error for the - 11/10

OpenStudy (anonymous):

no problem, ggodluck.

OpenStudy (anonymous):

does number 2 look right? 190 as the answer? for some reason i felt like i should have ponts x= (point,point) x=(point,point)

OpenStudy (anonymous):

*points

OpenStudy (anonymous):

no, 190 is correct, the problem is just asking what will be the value of f(x) when you substitute 8 to all the x's

OpenStudy (anonymous):

great! Thanks. life saver!

OpenStudy (anonymous):

you're welcome.